Circular No. 6346

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996L IN ESO 266-G10 R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of a supernova (mag ∼ 18.5), found on a U.K. Schmidt telescope R plate taken on Mar. 19.6 UT by K. S. Russell and J. M. James. SN 1996L is located 20′′.5 east and 2′′.2 north of the center of the host galaxy, at α = 11^{h}38^{m}37^{s}.80, δ = −43^{o}25′13′′.8 (equinox 2000.0; uncertainty 0′′.5 in each coordinate). A nearby star of mag ∼ 17.5 has position end figures 37^{s}.90, 6′′.1. No image appears at the position of SN 1996L on several surveys of various colors, the latest being in 1991.

S. Benetti, European Southern Observatory (ESO); P. Barthel, Kapteyn Astronomical Institute, Groningen; and W. de Vries, Space Telescope Science Institute, write: “Inspection of a preliminarily-reduced CCD spectrum (range 375.7–695.2 nm, resolution 1.5 nm) obtained on Mar. 21.15 UT with the ESO 3.6-m telescope (+ EFOSC1), confirms this object as a type-II peculiar supernova (IIn). Strong Hδ, Hγ, Hβ, and Hα emission lines are superimposed on a blue continuum (black-body temperature ∼ 11 500 K). The Balmer lines show a complex structure consisting of a narrow emission (FWHM ∼ 900 km/s), on top of a broader one (FWHM ∼ 4300 km/s). The He I (587.6 nm) line is also present in emission, with a FWHM of ∼ 2000 km/s. The recession velocity of the parent galaxy, as measured from emission lines of an H II region west of the supernova, is 9900 km/s.”

COMET C/1996 E1 (NEAT)

Improved parabolic orbital elements from MPEC 1996-F04:

T = 1996 July 27.3940 TT  ω = 81.^{o}3924
Ω = 149.8480 }2000.0
q = 1.352968 AU  i = 114.3714
1996 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Mar. 18 8^{h}53.^{m}69 +20^{o}41.^{′}0 1.486 2.278 132.^{o}3 18.^{o}9 16.4
23 8 40.52 +23 36.6 1.516 2.227 123.5 21.9 16.4
28 8 28.28 +26 16.7 1.556 2.176 115.1 24.5 16.3
Apr. 2 8 17.14 +28 40.7 1.604 2.126 107.2 26.7 16.3
7 8 07.21 +30 49.3 1.657 2.076  99.7 28.4 16.3
12 7 58.50 +32 44.0 1.715 2.026  92.7 29.6 16.2
17 7 51.03 +34 26.6 1.774 1.977  86.0 30.4 16.2
1996 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green